# UCINet0: A Machine Learning based Receiver for 5G NR PUCCH Format 0

### v1

@@ II-C1 Conventional Approaches
It is not clear how the thresholds in the above work and implementations extend to multiplexed UEs. Furthermore, fixed thresholds do not account for the received signal power variations that are caused due to variations in gains of hardware elements such as the LNA and due to varying locations of the multiplexed UEs. For example, a true transmission with large distance-based path loss could be misclassified as false if a fixed threshold is used.
Finding a one-size-fits-all threshold for various channel environments and multiplexed UEs is non-trivial. This has led us to explore an AI/ML-based data-driven approach in which a receiver could be designed directly from the raw frequency domain signals rather than correlation values. We apply similar techniques here. The IQ samples in the received resource block are analogous to pixel values in an image. Hence, models that work for image classification could potentially work here too.
II-C2 AI/ML Approaches
To our knowledge, there is limited work on applying AI/ML toward the decoding of Format 0 signals. One of the first models for Format 0 decoding was proposed in our previous work [6]. However, this was an initial effort that focused on training the model to predict 1 out of 4 possible $m_{cs}$ values in the specific case of 1 HARQ + 1 SR transmission. The architecture described in this paper extends our previous work to support all the possible UCI payloads shown in Table I and all the possible multiplexing combinations of UEs ($N_{UE}=0,1,\dots,12$). This generalized architecture is designed for potential field deployment in any 5G gNB.
III Machine Learning Receiver For PUCCH Format 0 Detection
@@
This section lays down the framework for posing the PUCCH Format 0 detection as an AI/ML classification problem. Specifically, we show how a single multi-label neural network classifier can serve as a generalized PUCCH Format 0 decoder.
III-A PUCCH Format 0 as an AI/ML Classification Problem
@@
Classification, a widespread use case of machine learning algorithms, is a supervised learning task that involves identifying to which class (label) a given data instance belongs. A predictive model, typically a Neural Network, is trained with a dataset that contains several input instances and the corresponding ground truth class labels. Labeled data helps us ”supervise” the NN in learning the correct input-output mapping.
@@
For PUCCH Format 0, the classification task is as follows: Given a sequence of received frequency domain Format 0 samples, predict the $N_{UE}$ phase rotation values $\alpha_{0},\alpha_{1}\dots\alpha_{N_{UE}-1}$ (see Eq. (3)) applied to the base sequence. The obtained $\alpha$ values can then be mapped back to the UCI-specific cyclic shift $m_{cs}$ for each UE by modulo 12 subtraction of the corresponding $m_{0}$ and $n_{cs}$ which are provided by the higher layers (L2). In the case where no user is transmitting on the PUCCH ($N_{UE}=0$), the NN prediction has to reflect this. In summary, the NN classifier has to predict either a single $\alpha$ value, multiple $\alpha$ values, or zero $\alpha$ values. Such a pattern of prediction is a typical use case of multi-label classification.
III-B Data Generation And Representation
@@
Dataset generation in AI/ML for wireless communication has a unique challenge not present in other domains - the non-availability of off-the-shelf benchmark datasets. One way of developing such datasets is through state-of-the-art simulation tools, such as the MATLAB 5G Toolbox, which can generate near-accurate datasets under various channel impairments. Simulated data is a good starting point for training neural networks in communication problems. However, we note that including field data, if available, gives us an insight into the generalization performance of AI/ML models across different distributions of data. The inclusion of field data also aligns with 3GPP requirements for 5G Release 18 [15]. For this paper, we use a combination of the above two approaches for data generation. The datasets used in this paper are publicly available at  [16].
III-B1 Simulated Datasets
@@
Using the MATLAB 5G Toolbox, we generate the received waveforms containing the PUCCH Format 0 signals. These generated waveforms include fading channel impairments and Additive White Gaussian Noise (AWGN). For various SNR values (0, 55, $10$, $15$, and $20$ dB), we generate PUCCH signals transmitted over a TDL-C Channel and store the noisy received samples. These samples are the input to the NN. For each input, the corresponding output label is the applied phase rotation $\alpha$ for each of the $N_{UE}$ UEs in multi-hot encoded format.
@@
The pseudocode for generating PUCCH Format 0 datasets in MATLAB is shown in Algorithm 1. We first place PUCCH Format 0 signals on all 14 symbols in a slot and 12 resource blocks in each symbol spaced 20 resource blocks apart. This results in 168 allocations per resource grid. We generate 1000 iterations of such resource grids. These 1000 grids are generated for each value of $N_{UE}$ in the range 0 to 12. The process is further repeated for the five SNR values.
@@
foreach $\operatorname{N_{UE}}$, $\operatorname{SNR}$, $\operatorname{Iter}$ do
@@
         Generate Format 0 signal and place it in the resource grid
@@
      Transmit over channel
@@
Algorithm 1 Training dataset generation
III-B2 Real-Time Over-The-Air Datasets

### v2

@@ II-D CONVENTIONAL RECEIVER METHODS FOR PUCCH FORMAT 0
The common metrics used in the above works are false and missed detections, where false detection refers to the detection of UCI when no UE has transmitted, and missed detection refers to the wrong detection of UCI. We note that the above works identify false transmissions by comparing the correlated peak with a fixed threshold. These fixed thresholds are specific to a certain scenario and are derived from lengthy Monte Carlo simulations. Another significant element that is not considered in [11, 12, 13] is the multiplexing of multiple UEs in the same resources. This is complicated by the fact that each UE can have different SNRs. It is not clear how the thresholds in the above work extend to multiplexed UEs. Furthermore, fixed thresholds do not account for variations in the received signal power caused by variations in gains of hardware elements such as the LNA and due to varying locations of the multiplexed UEs. For example, a true transmission with large distance-based path loss could be misclassified as false if a fixed threshold is used.
II-E CONVENTIONAL RECEIVER USING DFT-BASED CORRELATION
As can be seen in Eq. (6), the optimal Maximum Likelihood detector is an integer programming optimization problem and does not have a closed-form solution. In particular, solving the optimization problem would require a brute-force search over cyclic shifts $\bar{\alpha}$, transmission indicator $\beta_{i}$, and channels $h_{m}(k)$, which is very computationally expensive. A suboptimal receiver is based on the correlation with the base sequence without optimizing over the $\beta_{i}$ and the channel coefficients.
Subsequent to the correlation, the appropriate number of peaks is chosen based on the constraints. In many receivers, including the 5G Testbed deployment at IIT Madras [6], the correlation is implemented by an equivalent DFT-based algorithm for decoding the UCI of up to 12 multiplexed UEs. In this paper, we use this method as a baseline for comparison with the performance of the AI/ML model. Comparison with this particular DFT-based method is adequate because, to the best of our knowledge, the other existing academic literature on conventional decoders for Format 0, addresses only the single user scenario, i.e., only one user is transmitting a PUCCH Format 0 signal in a single resource.
Figure 2: Illustration of the correlation of the received Format 0 signal with the known base sequence for (a) 1 UE transmission and (b) 3 UEs multiplexed transmission
The first step of the DFT-based correlation is to multiply the received signal by the conjugate of the base sequence to obtain
$\tilde{y}(k)=y(k).\bar{r}^{*}_{u,v}(k)\stackrel{{\scriptstyle(a)}}{{=}}\sum_{m=0}^{N_{UE}-1}\beta_{m}h_{m}(k)e^{j\alpha_{m}k},k=0\ldots 11,$
where $(a)$ follows from the fact that $\|\bar{r}_{u,v}(k)\|^{2}=1$. Note that the noise $w(k)$ is omitted for the sake of being concise. Since the signal spans only one RB (for a $30$ kHz sub carrier spacing, this corresponds to $360$ KHz), a flat fading channel is assumed, ı.e., $h_{m}(k)=h_{m}$ for $k=0,1,\dots,11$. Hence, we obtain
$\tilde{y}(k)=\sum_{m=0}^{N_{UE}-1}\beta_{m}h_{m}e^{j\alpha_{m}k},~~k=0,1\ldots,11.$
We observe that $\tilde{y}(k)$ is a sum of complex exponentials with unknown weights, and the goal is to obtain the frequencies $\alpha_{m}$ of the different complex exponentials. A 12 point DFT is then applied to $\tilde{y}(k)$ which will peak at the frequencies that are present, i.e.,
$\hat{y}(k)=\text{DFT}(\tilde{y}(k)).$
Based on the constraints sets, the number of users, the location of the peaks of $\hat{y}(k)$ correspond to different $\alpha_{m}$ values. This is illustrated in Figure 2(a) with a correlation peak at $\text{index}=4$. The correlation index corresponding to the maximum correlation amplitude is taken as the estimated value $\alpha_{m}$. Recall from Eq. (2) that $\alpha$ is a combination of $m_{0}$, $m_{cs}$ and $n_{cs}$. Therefore, after subtracting $m_{0}+n_{cs}$ from $\alpha$ (subtraction is modulo 12), the cyclic shift specific to UCI $m_{cs}$ remains. In scenarios where multiple users ($N_{UE}$ of them) transmit PUCCH Format 0 signals at the same time-frequency location, the DFT-based correlation algorithm selects the top $N_{UE}$ peaks from the DFT output and uses them to determine $m_{cs}$ values for each of the $N_{UE}$ multiplexed UEs. Figure 2(b) shows an example in which 3 UEs are multiplexed. In this case, we select the top 3 peaks (at $\text{index}=4,6,\text{and }8$) as $\alpha_{m}$ values.
Observe that the above method is suboptimal, since we do not jointly optimize over $\beta_{m}$, and the constraint set. Also, the fading is assumed to be flat, which might not be the case with high mobility. However, in real deployments, the above technique is a very practical way of achieving correlation and provides good performance with reasonable implementation complexity.
III MACHINE LEARNING BASED RECEIVER FOR PUCCH FORMAT 0 DETECTION
@@
This section lays out the motivation and framework for posing the decoding of PUCCH Format 0 signals as an AI/ML classification problem. Specifically, we show how a single multi-label neural network classifier can serve as a generalized decoder for PUCCH Format 0.
The focus of this paper is only on Format 0 since its encoding of information in the phase of a waveform makes it suitable for formulation as an AI/ML classification problem i.e., the prediction of UCI content from a discrete set of phase rotations applied to a base sequence. It should be noted that since Format 0 is sequence based, its decoding does not require any information about the channel. In contrast, other 5G NR PUCCH formats use QPSK modulation of the UCI and require reference signals for channel estimation and equalization. The lengths of the time and frequency allocations are different and variable across the different formats. Other formats also involve scrambling, interleaving, rate matching and error correction mechanisms like Polar Coding and Reed Muller Coding. While it may be possible to bunch Formats 1 to 4 in a multi-task ML problem, Format 0 should be handled separately because of the significant differences in encoding and signal generation.
III-A MOTIVATION FOR THE APPLICATION OF AI/ML
Accurate decoding of Format 0 is important because of the following reasons:
  • Though there are 5 PUCCH formats, in the initial attach procedure, Format 0 is the most feasible Format that can be used to acknowledge and request resources for critical messages to set up a sustained communication link between a UE and a gNB.
  • Missed detection of a PUCCH Format 0 signal leads to the following scenarios:
    • A UE has decoded the gNB’s downlink data and acknowledged it with a HARQ transmission (on PUCCH Format 0 in the uplink direction). If this acknowledgement is missed at the gNB, the whole downlink message will be retransmitted, causing a significant latency in the message delivery and wastage of radio resources.
    • Another scenario is when a UE’s SR is missed by the gNB. In this case, the UE does not receive an uplink grant and the wait time for the successive grant is usually on the order of 10s of milliseconds. This causes the end-to-end latency of the system to increase.
  • False detection of a PUCCH Format 0 signal could result in unnecessary uplink grants and, as a consequence, radio resources being blocked.
The conventional model described above is based on correlating the received signal with all possible cyclically shifted versions of the base sequence and identifying valid transmissions (This can also be equivalently implemented by the DFT-based approach described above). This is the most common form of implementation of a Maximum Likelihood receiver for Format 0. We encounter the following two issues in designing such a decoder:
  • In a given time-frequency resource, the number of UEs scheduled to be multiplexed (by Layer 2) is merely an upper bound i.e., if $N_{UE}$ is the scheduled number of UEs, only $\tilde{N}_{UE}(\leq N_{UE})$ UEs among them might actually transmit. The difference between $N_{UE}$ and $\tilde{N}_{UE}$ is usually a consequence of one of the following (1) A UE does not transmit any SR despite a PUCCH Format 0 allocation, if no uplink grants are required because of a lack of data in its buffer (2) A UE does not transmit feedback (on HARQ resources) if it does not detect the prior downlink message. Since exact information about $\tilde{N}_{UE}$ is not known at the receiver, the receiver may falsely detect some transmissions. To mitigate this issue, a threshold could be used to invalidate false transmissions.
  • Fixed thresholds are hard to design because:
    • Based on the proximity to the base station, the correlation peak corresponding to each UE will vary.
    • In low SNR scenarios, true peaks are not easily distinguishable from noise.
    • The noise floor of the system can vary with temperature, front-end receiver characteristics, and other hardware-related parameters.
The above issues have led us to explore an AI/ML based approach. The data-driven nature of such an approach means that it can learn the hidden patterns from the complex IQ samples extracted directly from the resource grid.
TABLE III: Differences between prior work [5] and the UCINet0 architecture proposed in this work
Features Our prior work[5] This work (UCINet0)
Number of multiplexed UEs 1 Up to 12 (Max allowed by the spec)
Neural Network Input Rx IQ samples Rx IQ samples + metadata (number of scheduled UEs)
Inclusion of $\Delta$ - the mismatch between the actual number of multiplexed UEs and the upper bound indicated by the L2 scheduler No Yes
Inclusion of Constraint set mask for inference. See Section VI-B No Yes
Inclusion of multiple Doppler frequency shifts in the dataset No Yes
Neural Network output UCI specific cyclic shift ($m_{cs}$) Applied cyclic shift to the base sequence ($\alpha$)
Neural Network output size 4 12
Type of classification Multi-class (single UE) Multi-label (Multiple multiplexed UEs)
Number of combinations of UCI 1 (Only 1 SR + 1 HARQ) 5 (1 SR, 1 HARQ, 2 HARQ, 1 SR+ 1 HARQ, 2 HARQ + SR)
Identification of false transmissions No Yes
Insights into the interpretability of the model No Yes
To our knowledge, there is limited work on applying AI/ML toward the decoding of Format 0 signals. One of the models for Format 0 decoding was proposed in our previous work [5]. However, this was an initial effort that focused on training the model to predict 1 out of 4 possible $m_{cs}$ values in the specific case of 1 HARQ + 1 SR transmission. The UCINet0 architecture described in this paper extends our previous work to support all the possible UCI payloads shown in Table I and all the possible multiplexing combinations of UEs ($N_{UE}=0,1,\dots,12$). Table III highlights the key differences between our prior work and the generalized architecture described in this paper.
III-B PUCCH FORMAT 0 AS AN ML CLASSIFICATION PROBLEM
@@
Classification, a widespread use case of machine learning algorithms, is a supervised learning task that involves identifying the class (label) to which a given data instance belongs. A predictive model, typically a Neural Network (NN), is trained with a dataset that contains several input instances and the corresponding ground truth class labels. Labeled data helps us “supervise” the NN in learning the correct input-output mapping.
@@
Objective of the PUCCH Format 0 Classification Task: Given a sequence of received frequency domain Format 0 samples, predict the $\tilde{N}_{UE}$ phase rotation values $\alpha_{0},\alpha_{1}\dots\alpha_{\tilde{N}_{UE}-1}$ (see Eq. (3)) applied to the base sequence. The obtained $\alpha$ values can then be mapped back to the UCI-specific cyclic shift $m_{cs}$ for each UE by modulo 12 subtraction of the corresponding $m_{0}$ and $n_{cs}$, which are provided by the higher layers (L2). In the case where no user is transmitting on the PUCCH ($\tilde{N}_{UE}=0$), the NN prediction has to reflect this. In summary, the NN classifier has to predict either a single $\alpha$ value, multiple $\alpha$ values, or zero $\alpha$ values. Such a pattern of prediction is a typical use case of multi-label classification.
IV NEURAL NETWORK ARCHITECTURE
Given a problem statement, there is no formula for knowing apriori the best neural network architecture. For most problems, initial architectures are determined based on prior experience and domain knowledge. These architectures are then tuned by experiment. Figure 3 shows the UCINet0 architecture that showed the best performance in our experiments. In subsequent sections, we also comment on the trade-off between high complexity and performance.
Owing to the relatively small dimension of the input, our proposed UCINet0 is a Fully Connected Neural Network (FCN). The choice of an FCN also stems from the fact that the phase rotation, containing the UCI, is applied to all elements of the input vector rather than a localized portion (which is what CNNs would be good at extracting). Architectures such as RNNs, LSTM and transformers are not applicable in this case, since we are not dealing with sequential or time series data and there is no need to learn and retain long-term dependencies across the sequence i.e., there is no memory involved.
We also experimented with Complex Valued Neural Networks (CVNNs) in which the model inputs, outputs, and weights corresponding to each neuron are complex. Our ablation studies have shown that the real valued FCNs perform much better than their complex valued counterparts. We concluded that, given the small input size, the relationship between the real and imaginary parts of the input could just as easily be expressed by an equivalent FCN with twice the neurons in the input. However, CVNNs remain an interesting option for future applications of AI/ML in communications.
Figure 3: UCINet0 Architecture for PUCCH Format 0 decoding with 24 neurons in the input layer, 256+1 neurons in the second layer, 256 neurons in the third layer, and 12 neurons in the output layer.
IV-A ARCHITECTURE DESCRIPTION
The UCINet0 architecture, as shown in Figure. 3 has the following characteristics:
IV-A1 MODEL INPUT
The input to the NN has two components. The first component is the received PUCCH Format 0 signal. Since the PUCCH Format 0 signals occupy a single resource block, there are $N_{sc}^{RB}=12$ received complex samples denoted as $\overline{y}\in\mathbb{C}^{N_{sc}^{RB}\times 1}$. We split and concatenate the real and imaginary parts of $\overline{y}$ to obtain $\overline{Y}\in\mathbb{R}^{2N_{sc}^{RB}\times 1}$ as shown below. In this paper, $\overline{Y}$ constitutes the input layer of UCINet0.
$\begin{split}\overline{Y}&=[\text{Re}(\overline{y}),\text{ Im}(\overline{y})]\\ &=[\text{Re}(y(0)),\text{ Re}(y(1)),\dots\text{Re}(y(N_{sc}^{RB}-1)),\\ &\hskip 17.07164pt\text{Im}(y(0)),\text{ Im}(y(1)),\dots\text{Im}(y(N_{sc}^{RB}-1))]^{T}.\end{split}$ (7)
The second component of the input, which we feed to the first dense layer, is a metadata to further aid the NN. The metadata indicates the number of UEs scheduled by L2 ($N_{UE}$) in a given time-frequency resource. Even in practical 5G systems, the metadata $N_{UE}$ is known at the L1 (through FAPI protocol). From an AI/ML perspective, it can aid the model to learn information about the number of scheduled users.
However, as mentioned in the previous section, we note that the metadata from L2 need not always match the actual number of transmitting UEs. Hence we defined the quantity $\Delta$ (see (5)) to model this mismatch. In this paper, we consider various cases of a maximum metadata offset $\Delta$, which is the difference between the actual number of transmitting UEs and the scheduled number of multiplexed UEs. For example, $\Delta=0$ indicates that the number of UEs provided as the metadata input to the NN matches the number of UEs actually transmitting in the given allocation. $\Delta=2$ indicates that for each data instance, the number of UEs provided as the metadata input to the NN could be offset by any value among 0, 1 and 2, with 0 being the best case (no offset) and 2 being the worst. In this paper, during model training, we assume that $\Delta=2$ and we show inference results for $\Delta=0,2,4$.
IV-A2 HIDDEN LAYERS
We use 2 hidden layers of size 256. As stated above, the first hidden layer also takes the metadata as an input, in addition to the 256 neurons. Both hidden layers use a Rectified Linear Unit (ReLU) activation function.
IV-A3 MODEL OUTPUT
The output layer contains twelve neurons. Each output neuron acts as an indicator for the transmission of the base sequence cyclically shifted by $\alpha_{m}$ (Eq. (3)), where $m=0,1,2,\dots,11$.
The output layer uses a Sigmoid activation function. The use of the Sigmoid activation function is motivated by the fact that in a multi-label classifier, the sum of all the output probabilities need not be 1. Each output neuron can be thought of as a part of an independent binary classifier. In other words, since each output neuron corresponds to an $\alpha_{m}$ value, the probabilities of each output neuron indicate the confidence with which the NN detects the presence of the $m^{th}$ UE.
Figure 4: Model accuracy for different FCN architectures at various values of $\Delta$. Here, the number of FCN layers and the number of neurons of each architecture are presented on the x-axis, and $\Delta$ represents the maximum metadata offset between the true value and L2’s upper bound for the number of multiplexed UEs. We use these results to arrive at the final UCINet0 architecture presented in this paper.
Figure 5: Model test accuracy vs SNR for various values of $\Delta$. (a) No metadata given during training, (b) Metadata given during training at the input layer, (c) Metadata given during training at the first hidden layer, and (d) Metadata given during training at the second hidden layer.
Figure 6: Shapley Values for two example output classes (a) Class 1 (b) Class 5. The distribution of the Shapley values indicates that all input features (both received signal and the metadata) are necessary for the model’s prediction.
IV-B ABLATION STUDY: FCN ARCHITECTURES
We arrived at the final UCINet0 architecture of 2 layers and 256 neurons after evlauating the performance of other FCN architectures with a varying number of layers and neurons. Figure 4 shows the UCI decoding accuracies (defined in Section VII-B) obtained with 20 variations of FCN architectures. We considered combinations of $1$, 22, 33 and $4$ layer models with $64$, $128$, $256$, $512$ and $1024$ neurons in the hidden layers.
From Figure 4, we observe the following:
  • Model performance generally improves as we increase the number of neurons from $64$ to $1024$ for any number of layers and all values of $\Delta$.
  • For all values of $\Delta$, all of $1024$, $512$, and $256$ neurons result in almost the same accuracy values for a given number of hidden layers. For example, 2 $\text{ layers}\times 1024, 2$ $2\text{ layers}\times 51$2 and $2\text{ layers}\times 256$ result in similar accuracy values.
  • For all values of $\Delta$, we observe that the accuracy drops initially and then slowly increases as the number of layers increases. This trend becomes more explicit as we move from $1024$ to $128$ neuron models.
Hence, considering the balance between the number of layers, the number of neurons, and accuracies for various values of $\Delta$, the 256 neuron architecture was chosen. Though the $2\text{ layers}\times 256$ and $3\text{ layers}\times 256$ architectures exhibit similar performance, the former seemed to be the optimal choice because of its better performance when $\Delta=0$ and $2$, which are the most common metadata offset scenarios in practical systems.
Note that for this ablation study, in each architecture with more than 1 hidden layer, the penultimate hidden layer contains the metadata input as an additional neuron. For architectures with one hidden layer, the metadata is added to the input layer as an additional neuron. The motivation for the inclusion of such metadata is discussed in the subsequent section.
IV-C ABLATION STUDY: METADATA AS AN INPUT
To understand the need for metadata in training the neural network, we perform an ablation study of the model (1) with and without metadata, and (2) including metadata at various layers of the model.
Figure 5(a) shows the accuracy of the model across various SNRs when metadata is not fed as input to the neural network. Figure 5(b)5(d) show the accuracy with metadata fed in the input layer, the first dense layer, and the second dense layer, respectively. From Figure 5, it is clear that the inclusion of metadata provides significant gains. It can be observed that the exclusion of metadata leads to a $2\%$ accuracy drop at higher SNRs and $4\%$ accuracy drop at lower SNRs. We observe that the inclusion of metadata at the first hidden layer provides the best accuracy gain. This is reflected very prominently at lower SNRs. Consequently, we reflect this in our UCINet0 model architecture, such that the first hidden layer contains $256+1$ neurons as shown in Figure 3.)
IV-D SHAPLEY VALUES
To better understand the impact of various features of the input on the model prediction, we also look at the Shapley values, which explain the contribution of individual features towards a prediction. Figure 6, generated using the SHAP toolbox in Python shows the mean absolute SHAP values for each of the input features, for two example classes ($\alpha=1$ and $\alpha=5$). From the figure, we understand the following: (1) the metadata input (ch24) consistently has a high Shapley value, further emphasizing its importance in the model training; (2) among the other input features, i.e., the received PUCCH Format 0 signal, there is no dominant feature. Consider the feature ch01 as an example. In Figure 6(b), it has a low SHAP value, indicating its lower contribution for the prediction of Class $5$. However, the same feature has a high SHAP value in the prediction of Class $1$ as shown in Figure 6(a). Similar observations can be drawn from other features for different classes, indicating that the model requires all the features to make the best prediction. This conclusion validates our choice of an FCN architecture over CNN.
V DATA GENERATION
@@
Dataset generation in AI/ML for wireless communication has a unique challenge not present in other domains - the non-availability of off-the-shelf benchmark datasets. One way of developing such datasets is through state-of-the-art simulation tools, such as the MATLAB 5G Toolbox, which can generate near-accurate datasets under various channel impairments. Simulated data is a good starting point for training neural networks in communication problems. However, we note that including field data, if available, gives us an insight into the generalization performance of AI/ML models across different distributions of data. The inclusion of field data also aligns with 3GPP requirements for 5G Release 19 [14]. For this paper, we use a combination of the above two approaches for data generation. The datasets used in this paper are publicly available at [15].
V-A SIMULATED DATASETS
@@
Using the MATLAB 5G Toolbox, we generate received waveforms containing PUCCH Format 0 signals. These generated waveforms include fading channel impairments and Additive White Gaussian Noise (AWGN).
The fading channels include Doppler effects corresponding to frequency shifts ($f_{d}$) in the range of 0 to $2000$ Hz, which map to various UE speeds in the range of 0 (static) to $600$ kmph (high speed) for a carrier frequency of $3.5$ GHz. These Doppler frequency shifts allow for the simulation of fast-fading channels. For various SNR values in the range of 0 to $20$ dB, we generate PUCCH signals transmitted over various TDL channels (TDLA30, TDLA300, TDLC30 and TDL300) and store the noisy received samples. These samples are the input to the NN. For each input, the corresponding output label is the applied phase rotation value for each of the $N_{UE}$ UEs in a multi-hot encoded format. The pseudocode for generating PUCCH Format 0 datasets in MATLAB is shown in Algorithm 1.
@@
We first place PUCCH Format 0 signals on all 14 OFDM symbols in a slot and 12 resource blocks in each symbol, spaced 20 resource blocks apart. This results in 168 allocations per resource grid. We generate 1000 iterations (indicated by iter) of such resource grids. These 1000 grids are generated for each value of $N_{UE}$ in the range 0 to 12. The process is further repeated for the five SNR values. The size of the simulated dataset thus generated is $168000$ per $N_{UE}$ per SNR.
$\operatorname{\Delta}=\{0,2,4\}$
@@
foreach $\operatorname{\Delta}$, $\operatorname{N_{UE}}$, $\operatorname{SNR}$, $\operatorname{fd}$, $\operatorname{Iter}$ do
@@
         Generate a Format 0 signal and place it in the resource grid
@@
      Transmit over the channel
@@
Algorithm 1 Simulated dataset generation
V-B REAL-TIME OVER-THE-AIR DATASETS (LAB SETUP)